A triangular prism has bases that each have an approximate area of 173 square centimeters. The triangles measure 20 centimeters on each side and the prism is 4 centimeters high. What is the surface area.
1
Expert's answer
2020-04-17T18:34:33-0400

"Let's\;calculate\;area\;of\;lateral\;faces.\\S_{l.f.}=P\times h=(20+20+20)\times4=240\;cm^2\\Next,\\S_{bases}=173\times2=346\;cm^2\\Total\;Area\;=\;346\;+\;240\;=\;586\;cm^2\\"

Answer: 586cm2.